### v1

@@ 3.1 Some similarities
Two bounded curves $g_{1}$ and $g_{2}$ intersect if by traversing $g_{1}$ from one of its endpoints to the other endpoint, it crosses $g_{2}$ and switches from one side of $g_{2}$ to the other side [7, 23]. We say that $g_{1}$ and $g_{2}$ are non-intersecting if they do not intersect. Two non-intersecting curves can share an endpoint or can “touch” each other. Note that if $g_{1}$ and $g_{2}$ are geodesics they can intersect at most once [7].
It is well known that given two segments ${ab}$ and ${pq}$ that intersect transversely in the plane, any disk through $a$ and $b$ contains at least one endpoint of ${pq}$, or any disk through $p$ and $q$ contains at least one endpoint of ${ab}$, see, e.g., [18]. This result translates to the geodesic setting as shown below. We begin with a property of geodesic disks.
@@
Let $a,b,p,q$ be four distinct points in general position in $P$ such that $g(p,q)$ and $g(a,b)$ intersect. Then every geodesic disk with $p$ and $q$ on its boundary contains at least one endpoint of $g(a,b)$, or every geodesic disk with $a,b$ on its boundary contains at least one endpoint of $g(p,q)$.
@@
First, suppose that the bisectors $b(a,b)$ and $b(p,q)$ intersect, see Figure 3(a), and let $x$ be a point in their intersection. Then we have that $|g(x,a)|=|g(x,b)|$ and $|g(x,p)|=|g(x,q)|$. We can assume, without loss of generality, that $|g(x,p)|<|g(x,a)|$ since the points in $S$ are in general position, and $a,b,p,q$ are not co-circular. The case where $|g(x,p)|>|g(x,a)|$ is symmetric.
@@
When $|g(x,p)|<|g(x,a)|$, the geodesic disk through $a$ and $b$ with center $x$ contains $p$ and $q$. Since $g(a,b)$ and $g(p,q)$ intersect, $p$ and $q$ are on different sides of the extension path $\ell(a,b)$. Hence, by Lemma 1, if we move the center of this disk along the bisector $b(a,b)$ in one direction while keeping $a,b$ on its boundary, the geodesic disk will always contain $p$ in its interior, and if we move the center in the other direction, the geodesic disk will always contain $q$.
@@
Now, suppose that the bisectors $b(a,b)$ and $b(p,q)$ do not intersect, see Figure 3(b). Hence, all points of the bisector $b(a,b)$ are closer to one endpoint of $g(p,q)$ than to the other one. Assume, without loss of generality, that this point is $p$, i.e., $|g(x^{\prime},p)|\leq|g(x^{\prime},q)|,\ \forall x^{\prime}\in b(a,b)$. Similarly, all points of the bisector $b(p,q)$ are closer to one endpoint of $g(a,b)$ than to the other one. Assume, without loss of generality, that this point is $b$.
(a) Bisectors $b(a,b)$ and $b(p,q)$ intersect.
(b) Bisectors $b(a,b)$ and $b(p,q)$ do not intersect.
Figure 3: Any geodesic disk through $a$ and $b$ contains at least one endpoint of $g(p,q)$.
@@
The bisector $b(a,b)$ intersects the boundary of the polygon. Also, the extension path $\ell(a,b)$ divides the polygon $P$ into two half-polygons, one containing $p$ and the other containing $q$. Let $x$ be the intersection of $b(a,b)$ with the polygon boundary that is contained in the same half-polygon as $q$. Similarly, the extension path $\ell(p,q)$ divides the polygon into two half-polygons, and let $y$ be the intersection of $b(p,q)$ with the polygon contained in the same half-polygon as $a$.
@@
By definition, we have that $|g(x,a)|=|g(x,b)|$, and $|g(x,p)|<|g(x,q)|$. If $|g(x,p)|\leq|g(x,b)|$, the geodesic disk with center at $x$ and radius $|g(x,b)|$ contains $a,b$ and $p$. By Lemma 1, this means that every geodesic disk through $a,b$ contains $p$.
@@
The only remaining case to consider is if $|g(x,p)|>|g(x,b)|$. Since the bisectors do not intersect, we have that $g(x,b)$ intersects $g(y,p)$. Let $z$ be a point in this intersection. By definition, we have $|g(x,b)|=|g(x,z)|+|g(z,b)|$ and $|g(y,p)|=|g(y,z)|+|g(z,p)|$. By the triangle inequality, we have $|g(x,p)|\leq|g(x,z)|+|g(z,p)|$, which with our assumption that $|g(x,p)|>|g(x,b)|$ implies that $|g(z,b)|<|g(z,p)|$. However, this means that $|g(y,b)|\leq|g(y,z)|+|g(z,b)|<|g(y,z)|+|g(z,p)|=|g(y,p)|$. Therefore, when $|g(x,p)|>|g(x,b)|$, we have $|g(y,b)|<|g(y,p)|$. By Lemma 1, this means that every geodesic disk through $p,q$ contains $b$. ∎
@@
Let $a,b,p,q$ be four distinct points in general position in $P$ such that $g(p,q)$ and $g(a,b)$ intersect. Then every geodesic disk with $p$ and $q$ on its boundary does not contain at least one endpoint of $g(a,b)$, or every geodesic disk with $a,b$ on its boundary does not contain at least one endpoint of $g(p,q)$.
@@
Suppose that the bisectors $b(a,b)$ and $b(p,q)$ intersect, see Figure 3(a), and let $x$ be a point in their intersection. Then we have that $|g(x,a)|=|g(x,b)|$ and $|g(x,p)|=|g(x,q)|$. We can assume, without loss of generality, that $|g(x,a)|<|g(x,p)|$ since the points in $S$ are in general position, and $a,b,p,q$ are not co-circular. The case where $|g(x,a)|>|g(x,p)|$ is symmetric.
@@
When $|g(x,a)|<|g(x,p)|$, the geodesic disk through $a$ and $b$ with center $x$ contains neither $p$ nor $q$. Since $g(a,b)$ and $g(p,q)$ intersect, $p$ and $q$ are on different sides of the extension path $\ell(a,b)$. Note that if we exchange the roles of $p$ by $q$, and $D$ by $D^{\prime}$ in Lemma 1, we have that $D^{\prime}\cap P^{-}(p,q)\subset D$. Hence, by Lemma 1, if we move the center of this disk along the bisector $b(a,b)$ in one direction while keeping $a,b$ on its boundary, the geodesic disk will never contain $p$, and if we move the center in the other direction, the geodesic disk will never contain $q$.
@@
Now, suppose that the bisectors $b(a,b)$ and $b(p,q)$ do not intersect, see Figure 3(b). Hence, all points of the bisector $b(a,b)$ are closer to one endpoint of $g(p,q)$ than to the other one. Assume, without loss of generality, that this point is $p$, i.e., $|g(x^{\prime},p)|\leq|g(x^{\prime},q)|,\ \forall x^{\prime}\in b(a,b)$. Similarly, all points of the bisector $b(p,q)$ are closer to one endpoint of $g(a,b)$ than to the other one. Assume, without loss of generality, that this point is $b$.
@@
The bisector $b(a,b)$ intersects the boundary of the polygon. Also, the extension path $\ell(a,b)$ divides the polygon $P$ into two half-polygons, one containing $p$ and the other containing $q$. Let $x$ be the intersection of $b(a,b)$ with the polygon boundary that is contained in the same half-polygon as $q$. Similarly, the extension path $\ell(p,q)$ divides the polygon into two half-polygons, and let $y$ be the intersection of $b(p,q)$ with the polygon contained in the same half-polygon as $a$.
@@
By definition, we have that $|g(x,a)|=|g(x,b)|$, and $|g(x,p)|<|g(x,q)|$. If $|g(x,b)|<|g(x,q)|$, the geodesic disk with center at $x$ through $a,b$ does not contain $q$. By Lemma 1, this means that every geodesic disk through $a,b$ does not contain $q$.
@@
The only remaining case to consider is when $|g(x,b)|\geq|g(x,q)|$. By the fact that $a,b,p,q$ are not co-circular, we have that in fact $\lvert g(x,b)\rvert>\lvert g(x,q)\rvert$. Since the bisectors do not intersect, we have that $g(x,q)$ intersects $g(y,b)$. Let $z$ be a point in this intersection. By definition, we have $|g(x,q)|=|g(x,z)|+|g(z,q)|$ and $|g(y,b)|=|g(y,z)|+|g(z,b)|$. By the triangle inequality, we have $|g(x,b)|\leq|g(x,z)|+|g(z,b)|$, which with our assumption that $|g(x,b)|>|g(x,q)|$ implies that $|g(z,b)|>|g(z,p)|$. However, this means that $|g(y,q)|<|g(y,z)|+|g(z,q)|\leq|g(y,z)|+|g(z,b)|=|g(y,b)|$. Therefore, when $|g(x,b)|\geq|g(x,q)|$, we have $|g(y,p)|<|g(y,b)|$. By Lemma 1, this means that every geodesic disk through $p,q$ does not contain $b$. ∎
@@
Let $\triangle(u,v,w)$ be a geodesic triangle, such that $v$ is a convex vertex of $\triangledown$$\triangledown(u,v,w)$ and $\angle uvw\geq\frac{\pi}{3}$. Then $|g(u,w)|\geq\min\{|g(u,v)|,|g(v,w)|\}$.
@@
Figure 4: The angle $\angle uvw$ is greater than $\frac{\pi}{3}$, and $|vq|<|vp|$.
Since $\triangle(v,p,q)$ is a Euclidean triangle, we know that $|pq|\geq\min\{|vp|,|vq|\}$. We can assume, without loss of generality, that $|pq|\geq|vq|$. Let $x$ be the vertex that is the endpoint of the first edge of $g(v,u)$. By construction $|vq|=|vx|+|xq|$. Note that $|g(p,q)|\geq|pq|\geq|vq|=|vx|+|xq|$ (1). By the triangle inequality, $|g(x,u^{\prime})|\leq|xq|+|g(q,u^{\prime})|$. This implies that $|g(x,u^{\prime})|-|xq|\leq|g(q,u^{\prime})|$ (2). Putting it all together, we have
$|g(v,u)|$ $=|g(u,u^{\prime})|+|g(v,u^{\prime})|\mbox{ by definition}$
$=|g(u,u^{\prime})|+|vx|+|g(x,u^{\prime})|\mbox{ by definition}$
$=|g(u,u^{\prime})|+|vx|+|g(x,u^{\prime})|-|xq|+|xq|$
[…]
$\leq|g(p,u)|$
$\leq|g(w,u)|$
We conclude that $|g(u,w)|\geq\min\{|g(u,v)|,|g(v,w)|\}$. ∎
@@
Let $\triangle(u,v,w)$ be a geodesic triangle, such that $\angle uvw\leq\frac{\pi}{3}$. Then we have that $|{uw}|\leq\max\{|g(u,v)|,|g(v,w)|\}$.
@@
Since the three points $u,v,w$ form a Euclidean triangle, we know that $|{uw}|\leq\max\{|{uv}|,|{vw}|\}$. We can assume, without loss of generality, that $|{uw}|\leq|{uv}|$. Since $|{uv}|\leq|g(u,v)|$, we thus have that $|{uw}|\leq|g(u,v)|\leq\max\{|g(u,v)|,|g(v,w)|\}$. ∎

### v2

@@ 3.1 Some similarities
Two bounded curves $g_{1}$ and $g_{2}$ intersect if by traversing $g_{1}$ from one of its endpoints to the other endpoint, it crosses $g_{2}$ and switches from one side of $g_{2}$ to the other side [7, 23]. We say that $g_{1}$ and $g_{2}$ are non-intersecting if they do not intersect. Two non-intersecting curves can share an endpoint or can “touch” each other. Note that if $g_{1}$ and $g_{2}$ are geodesics, they can intersect at most once [7], i.e., their intersection consists of at most one subpath of both geodesics.
It is well known that given two segments $uv$ and $pq$ that intersect transversely in the plane, any disk through $u$ and $v$ contains at least one endpoint of $pq$, or any disk through $p$ and $q$ contains at least one endpoint of $uv$, see, e.g., [18]. This result translates to the geodesic setting as shown below. We begin with a property of geodesic disks, refer to Figure 3.
Figure 3: The intersection of the disk through $p$ and $q$ with center at $c$ (in red) and $P^{+}(p,q)$ (in gray) is contained in the disk through $p$ and $q$ with center at $c^{\prime}$ (in blue).
@@
Let $u,v,p,q$ be four distinct points in general position in $P$ such that $g(p,q)$ and $g(u,v)$ intersect. Then every geodesic disk with $p$ and $q$ on its boundary contains at least one endpoint of $g(u,v)$, or every geodesic disk with $u,v$ on its boundary contains at least one endpoint of $g(p,q)$.
@@
First, suppose that the bisectors $b(u,v)$ and $b(p,q)$ intersect, see Figure 4(a), and let $x$ be a point in their intersection. Then we have that $\lvert g(x,u)\rvert=\lvert g(x,v)\rvert$ and $\lvert g(x,p)\rvert=\lvert g(x,q)\rvert$. We can assume, without loss of generality, that $\lvert g(x,p)\rvert<\lvert g(x,u)\rvert$ since the points in $S$ are in general position, and $u,v,p,q$ are not co-circular. The case where $\lvert g(x,p)\rvert>\lvert g(x,u)\rvert$ is symmetric.
@@
When $\lvert g(x,p)\rvert<\lvert g(x,u)\rvert$, the geodesic disk through $u$ and $v$ with center $x$ contains $p$ and $q$. Since $g(u,v)$ and $g(p,q)$ intersect, $p$ and $q$ are on different sides of the extension path $\ell(u,v)$. Hence, by Lemma 1, if we move the center of this disk along the bisector $b(u,v)$ in one direction while keeping $u,v$ on its boundary, the geodesic disk will always contain $p$ in its interior, and if we move the center in the other direction, the geodesic disk will always contain $q$.
@@
Now, suppose that the bisectors $b(u,v)$ and $b(p,q)$ do not intersect, see Figure 4(b). Hence, all points of the bisector $b(u,v)$ are closer to one endpoint of $g(p,q)$ than to the other one. Assume, without loss of generality, that this point is $p$, i.e., $\lvert g(x^{\prime},p)\rvert\leq\lvert g(x^{\prime},q)\rvert,\ \forall x^{\prime}\in b(u,v)$. Similarly, all points of the bisector $b(p,q)$ are closer to one endpoint of $g(u,v)$ than to the other one. Assume, without loss of generality, that this point is $v$.
(a) Bisectors $b(u,v)$ and $b(p,q)$ intersect.
(b) Bisectors $b(u,v)$ and $b(p,q)$ do not intersect.
Figure 4: Any geodesic disk through $u$ and $v$ contains at least one endpoint of $g(p,q)$.
@@
The bisector $b(u,v)$ intersects the boundary of the polygon. Also, the extension path $\ell(u,v)$ divides the polygon $P$ into two half-polygons, one containing $p$ and the other containing $q$. Let $x$ be the intersection of $b(u,v)$ with the polygon boundary that is contained in the same half-polygon as $q$. Similarly, the extension path $\ell(p,q)$ divides the polygon into two half-polygons, and let $y$ be the intersection of $b(p,q)$ with the polygon contained in the same half-polygon as $u$.
@@
By definition, we have that $\lvert g(x,u)\rvert=\lvert g(x,v)\rvert$, and $\lvert g(x,p)\rvert<\lvert g(x,q)\rvert$. If $\lvert g(x,p)\rvert\leq\lvert g(x,v)\rvert$, the geodesic disk with center at $x$ and radius $\lvert g(x,v)\rvert$ contains $u,v$ and $p$. By Lemma 1, this means that every geodesic disk through $u,v$ contains $p$.
@@
The only remaining case to consider is when $\lvert g(x,p)\rvert>\lvert g(x,v)\rvert$. Since the bisectors do not intersect, we have that $g(x,v)$ intersects $g(y,p)$. Let $z$ be a point in this intersection. By definition, we have $\lvert g(x,v)\rvert=\lvert g(x,z)\rvert+\lvert g(z,v)\rvert$ and $\lvert g(y,p)\rvert=\lvert g(y,z)\rvert+\lvert g(z,p)\rvert$. By the triangle inequality, we have $\lvert g(x,p)\rvert\leq\lvert g(x,z)\rvert+\lvert g(z,p)\rvert$, which with our assumption that $\lvert g(x,p)\rvert>\lvert g(x,v)\rvert$ implies that $\lvert g(z,v)\rvert<\lvert g(z,p)\rvert$. However, this means that $\lvert g(y,v)\rvert\leq\lvert g(y,z)\rvert+\lvert g(z,v)\rvert<\lvert g(y,z)\rvert+\lvert g(z,p)\rvert=\lvert g(y,p)\rvert$. Therefore, when $\lvert g(x,p)\rvert>\lvert g(x,v)\rvert$, we have $\lvert g(y,v)\rvert<\lvert g(y,p)\rvert$. By Lemma 1, this means that every geodesic disk through $p,q$ contains $v$. ∎
@@
Let $u,v,p,q$ be four distinct points in general position in $P$ such that $g(p,q)$ and $g(u,v)$ intersect. Then every geodesic disk with $p$ and $q$ on its boundary does not contain at least one endpoint of $g(u,v)$, or every geodesic disk with $u,v$ on its boundary does not contain at least one endpoint of $g(p,q)$.
@@
Suppose that the bisectors $b(u,v)$ and $b(p,q)$ intersect, see Figure 4(a), and let $x$ be a point in their intersection. Then we have that $\lvert g(x,u)\rvert=\lvert g(x,v)\rvert$ and $\lvert g(x,p)\rvert=\lvert g(x,q)\rvert$. We can assume, without loss of generality, that $\lvert g(x,u)\rvert<\lvert g(x,p)\rvert$ since the points in $S$ are in general position, and $u,v,p,q$ are not co-circular. The case where $\lvert g(x,u)\rvert>\lvert g(x,p)\rvert$ is symmetric.
@@
When $\lvert g(x,u)\rvert<\lvert g(x,p)\rvert$, the geodesic disk through $u$ and $v$ with center $x$ contains neither $p$ nor $q$. Since $g(u,v)$ and $g(p,q)$ intersect, $p$ and $q$ are on different sides of the extension path $\ell(u,v)$. Note that if we exchange the roles of $p$ by $q$, and $D$ by $D^{\prime}$ in Lemma 1, we have that $D^{\prime}\cap P^{-}(p,q)\subset D$. Hence, by Lemma 1, if we move the center of this disk along the bisector $b(u,v)$ in one direction while keeping $u,v$ on its boundary, the geodesic disk will never contain $p$, and if we move the center in the other direction, the geodesic disk will never contain $q$.
@@
Now, suppose that the bisectors $b(u,v)$ and $b(p,q)$ do not intersect, see Figure 4(b). Hence, all points of the bisector $b(u,v)$ are closer to one endpoint of $g(p,q)$ than to the other one. Assume, without loss of generality, that this point is $p$, i.e., $\lvert g(x^{\prime},p)\rvert\leq\lvert g(x^{\prime},q)\rvert,\ \forall x^{\prime}\in b(u,v)$. Similarly, all points of the bisector $b(p,q)$ are closer to one endpoint of $g(u,v)$ than to the other one. Assume, without loss of generality, that this point is $v$.
@@
The bisector $b(u,v)$ intersects the boundary of the polygon. Also, the extension path $\ell(u,v)$ divides the polygon $P$ into two half-polygons, one containing $p$ and the other containing $q$. Let $x$ be the intersection of $b(u,v)$ with the polygon boundary that is contained in the same half-polygon as $q$. Similarly, the extension path $\ell(p,q)$ divides the polygon into two half-polygons, and let $y$ be the intersection of $b(p,q)$ with the polygon contained in the same half-polygon as $u$.
@@
By definition, we have that $\lvert g(x,u)\rvert=\lvert g(x,v)\rvert$, and $\lvert g(x,p)\rvert<\lvert g(x,q)\rvert$. If $\lvert g(x,v)\rvert<\lvert g(x,q)\rvert$, the geodesic disk with center at $x$ through $u,v$ does not contain $q$. By Lemma 1, this means that every geodesic disk through $u,v$ does not contain $q$.
@@
The only remaining case to consider is when $\lvert g(x,v)\rvert\geq\lvert g(x,q)\rvert$. By the fact that $u,v,p,q$ are not co-circular, we have that in fact $\lvert g(x,v)\rvert>\lvert g(x,q)\rvert$. Since the bisectors do not intersect, we have that $g(x,q)$ intersects $g(y,v)$. Let $z$ be a point in this intersection. By definition, we have $\lvert g(x,q)\rvert=\lvert g(x,z)\rvert+\lvert g(z,q)\rvert$ and $\lvert g(y,v)\rvert=\lvert g(y,z)\rvert+\lvert g(z,v)\rvert$. By the triangle inequality, we have $\lvert g(x,v)\rvert\leq\lvert g(x,z)\rvert+\lvert g(z,v)\rvert$, which with our assumption that $\lvert g(x,v)\rvert>\lvert g(x,q)\rvert$ implies that $\lvert g(z,v)\rvert>\lvert g(z,p)\rvert$. However, this means that $\lvert g(y,q)\rvert<\lvert g(y,z)\rvert+\lvert g(z,q)\rvert\leq\lvert g(y,z)\rvert+\lvert g(z,v)\rvert=\lvert g(y,v)\rvert$. Therefore, when $\lvert g(x,v)\rvert\geq\lvert g(x,q)\rvert$, we have $\lvert g(y,p)\rvert<\lvert g(y,v)\rvert$. By Lemma 1, this means that every geodesic disk through $p,q$ does not contain $v$. ∎
@@
Let $\triangle(u,v,w)$ be a geodesic triangle, such that $v$ is a convex vertex of $\triangledown$$\triangledown(u,v,w)$ and $\angle uvw\geq\frac{\pi}{3}$. Then $\lvert g(u,w)\rvert\geq\min\{\lvert g(u,v)\rvert,\lvert g(v,w)\rvert\}$.
@@
Figure 5: The angle $\angle uvw$ is greater than $\frac{\pi}{3}$, and $\lvert vq\rvert<\lvert vp\rvert$.
Since $\triangle(v,p,q)$ is a Euclidean triangle, we know that $\lvert pq\rvert\geq\min\{\lvert vp\rvert,\lvert vq\rvert\}$. We can assume, without loss of generality, that $\lvert pq\rvert\geq\lvert vq\rvert$. Let $x$ be the vertex that is the endpoint of the first edge of $g(v,u)$. By construction $\lvert vq\rvert=\lvert vx\rvert+\lvert xq\rvert$. Note that $\lvert g(p,q)\rvert\geq\lvert pq\rvert\geq\lvert vq\rvert=\lvert vx\rvert+\lvert xq\rvert$ (1). By the triangle inequality, $\lvert g(x,u^{\prime})\rvert\leq\lvert xq\rvert+\lvert g(q,u^{\prime})\rvert$. This implies that $\lvert g(x,u^{\prime})\rvert-\lvert xq\rvert\leq\lvert g(q,u^{\prime})\rvert$ (2). Putting it all together, we have
$\lvert g(v,u)\rvert$ $=\lvert g(u,u^{\prime})\rvert+\lvert g(v,u^{\prime})\rvert\mbox{ by definition}$
$=\lvert g(u,u^{\prime})\rvert+\lvert vx\rvert+\lvert g(x,u^{\prime})\rvert\mbox{ by definition}$
$=\lvert g(u,u^{\prime})\rvert+\lvert vx\rvert+\lvert g(x,u^{\prime})\rvert-\lvert xq\rvert+\lvert xq\rvert$
$\leq\lvert g(u,u^{\prime})\rvert+\lvert vx\rvert+\lvert g(q,u^{\prime})\rvert+\lvert xq\rvert\mbox{ by (2)}$
$\leq\lvert g(q,u)\rvert+\lvert g(p,q)\rvert\mbox{ by (1)}$
$\leq\lvert g(p,u)\rvert$
$\leq\lvert g(w,u)\rvert$
We conclude that $\lvert g(u,w)\rvert\geq\min\{\lvert g(u,v)\rvert,\lvert g(v,w)\rvert\}$. ∎
The next lemma presents a result analogous to Lemma 4, but when the angle $\angle uvw$ is at most $\frac{\pi}{3}$.
@@
Let $\triangle(u,v,w)$ be a geodesic triangle, such that $\angle uvw\leq\frac{\pi}{3}$. Then we have that the edge $\lvert uw\rvert\leq\max\{\lvert g(u,v)\rvert,\lvert g(v,w)\rvert\}$.
@@
Since the three points $u,v,w$ form a Euclidean triangle, we know that $\lvert uw\lvert\leq\max\{\lvert uv\rvert,\lvert vw\lvert\}$. We can assume, without loss of generality, that $\lvert uw\lvert\leq\lvert uv\rvert$. Since $\lvert uv\rvert\leq\lvert g(u,v)\rvert$, we thus have that $\lvert uw\lvert\leq\lvert g(u,v)\rvert\leq\max\{\lvert g(u,v)\rvert,\lvert g(v,w)\rvert\}$. ∎